OpenStudy (anonymous):

Substitution Method 2x – y = -4 -3x + y = -9

OpenStudy (anonymous):

is x=13 and y=30

OpenStudy (anonymous):

i dont know i have to get answer show steps and check step

OpenStudy (anonymous):

sorry yes that is right but i need how you got that and the check step

OpenStudy (anonymous):

you need help?

OpenStudy (anonymous):

the equation is already set up for you

OpenStudy (anonymous):

in other words, you got your negative y and your positive y

OpenStudy (anonymous):

so you just add the two equations together

OpenStudy (anonymous):

it comes out to be -x=-13

OpenStudy (anonymous):

which simplifies to x=13

OpenStudy (anonymous):

to solve for y

OpenStudy (anonymous):

you need to see whats a common multiple for 2 and 3

OpenStudy (anonymous):

you also need one x to be positive while the other x is negative

OpenStudy (anonymous):

a common multiple would be 6

OpenStudy (anonymous):

so multiply 3 by 2 and 2 by 3

OpenStudy (anonymous):

and you get 6x plus -6x

OpenStudy (anonymous):

opps

OpenStudy (anonymous):

actually multiply the first equation by -3

OpenStudy (anonymous):

and the second equation by positive 2

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

you know what you could do

OpenStudy (anonymous):

use substitution

OpenStudy (anonymous):

in other words, plug in x=13 into both equations

OpenStudy (anonymous):

that will lead you to the conclusion that y=30

OpenStudy (anonymous):

you need any more explanations?

OpenStudy (anonymous):

I have taken Pre-Algebra, Algebra 1, and Algebra 2

OpenStudy (anonymous):

Although I'm not an official helper member for this website

OpenStudy (anonymous):

i dont get this stuff at all i can get the answers from a site but i cant explain how i got them no matter what i try to do thank you so muchi apprate it

OpenStudy (anonymous):

do you understand now?

OpenStudy (anonymous):

i got all kinds of problems with my math

OpenStudy (anonymous):

or do I need to break it down into simpler terms?

OpenStudy (anonymous):

someish

OpenStudy (anonymous):

k

OpenStudy (anonymous):

hmm...

OpenStudy (anonymous):

2x – y = -4 -3x + y = -9

OpenStudy (anonymous):

so

OpenStudy (anonymous):

you want to find x or y first?

OpenStudy (anonymous):

i dont know does it matter which one goes first just a question

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

but y is easier

OpenStudy (anonymous):

do you know why?

OpenStudy (anonymous):

hint hint: +,-.

OpenStudy (anonymous):

hint hint hint: canceling positives and negatives...

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

HINT HINT: lolz, If I have three apples and I took three away, how much do I have left?

OpenStudy (anonymous):

none

OpenStudy (anonymous):

there we go

OpenStudy (anonymous):

you understanding so far?

OpenStudy (anonymous):

yea its a start lol

OpenStudy (anonymous):

so all you have left is (2x=-4) +( -3x=-9)

OpenStudy (anonymous):

do you know how to add positives and negatives?

OpenStudy (anonymous):

its ok if you don't

OpenStudy (anonymous):

my calulator?

OpenStudy (anonymous):

besides the calculator... :D

OpenStudy (anonymous):

like lets see I have two copies of Skyrim and someone takes three of them away, how many do I have left?

OpenStudy (anonymous):

In other words, someone take two of the real Skyrim and took an imaginary one

OpenStudy (anonymous):

So we have -1 Skyrims now

OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

k, so whats -x=-13?

OpenStudy (anonymous):

a negative divided by a negative is...

OpenStudy (anonymous):

sorry puter is being slow not ignoring you

OpenStudy (anonymous):

its fine mon

OpenStudy (anonymous):

what makes a person content: A. Fifty piles of manure. B. A bill of 50,000 dollars. C. Cow manure. D. 5 trillion dollars

OpenStudy (anonymous):

this will lead you to the answer

OpenStudy (anonymous):

but i dont kkn ow what x is to divide -13 by

OpenStudy (anonymous):

What does this mean in mathematics, positive or negative?

OpenStudy (anonymous):

k

OpenStudy (anonymous):

a negative divided by a negative is always, always, ice cream cones always a positive

OpenStudy (anonymous):

since -x and -13 are both negatives, a negative divided by a negative yields a...

OpenStudy (anonymous):

I'll give you the first letter of the word

OpenStudy (anonymous):

P

OpenStudy (anonymous):

positive

OpenStudy (anonymous):

CORRECT!!!!!!!!!!!!!

OpenStudy (anonymous):

WE HAVE A WINNER!!!!!!!!!!!!!

OpenStudy (anonymous):

so x=13

OpenStudy (anonymous):

do you need further help to solve for y?

OpenStudy (anonymous):

hint hint: plug in the value for x(hint hint hint: 2(13)=26 and -3(13) 39

OpenStudy (anonymous):

I mean -39

OpenStudy (anonymous):

i think i understand how do i check it

OpenStudy (anonymous):

k

OpenStudy (anonymous):

do you know how to solve 26-y=-4?

OpenStudy (anonymous):

sorry i dont

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

so...

OpenStudy (anonymous):

you subtract 26 by both sides

OpenStudy (anonymous):

so you get -y=-30

OpenStudy (anonymous):

this remind you of Pre-Algebra at all?

OpenStudy (anonymous):

i have never done any algrbra class

OpenStudy (anonymous):

then why are you doing algebraic formulas?

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

and if you need help with basic algebraic formulas

OpenStudy (anonymous):

http://www.khanacademy.org/

OpenStudy (anonymous):

thats a very good website for learning the basics of Algebra

OpenStudy (anonymous):

and if that bores you too much

OpenStudy (anonymous):

just ask me for some Algebraic tips

OpenStudy (anonymous):

k?

OpenStudy (anonymous):

or ask people like eyust707

OpenStudy (anonymous):

or dpalnc